Wine Glass Inc. produces and sells wine glasses.  

Per the Standards, for the months of May, each wine glass takes 2 pounds of material at $2.50 per pound.

For the Month of May, Wine Glass Inc. planned on producing and selling 200 wine glasses.  They actually produced and sold 210 wine glasses, using 525 pounds of material costing $1,181.25 in actual total cost.

  1. What is the Materials Activity Variance for the Month of May
  2. What is the Materials Spending Variance for the Month of May
  3. What is the Price Variance:
  4. What is the Quantity Variance:
Answer #1
B Spending variance (AQ*AP) - (SQ*SP) AQ= ACTUAL QTY
(210*1181.25) - (200*1125) AP= ACTUAL PRICE
248062.5-225000 SQ= STND QTY
$23062.5 SP= STND PRICE
Favourable
C Price variance AQ-(AP-SP)
210*(1181.25-1125)
210*56.25
$11812.5
Favourable
D Quantity vaiance SP*(AQ-SQ)
1125*(210-200)
$11250
Favourable
****Material Activity Variance There is no activity variance in material variance, Hence, unable to answer the activity volume..

ANSWER IS SOLVED AS PER GIVEN QUESTION.
If you like the answer, kindly raise your UPVOTE button.
ALL THE BEST!!!!

if any doubts ask!!! #STAY HOME,STAY SAFE